# TEA: Test-time Energy Adaptation

### 24 Nov 2023 v1

@@ Abstract
Test-time adaptation (TTA) aims to improve model generalizability when test data diverges from training distribution, offering the distinct advantage of not requiring access to training data and processes, especially valuable in the context of large pre-trained models. However, current TTA methods fail to address the fundamental issue: covariate shift, i.e., the decreased generalizability can be attributed to the model’s reliance on the marginal distribution of the training data, which may impair model calibration and introduce confirmation bias. To address this, we propose a novel energy-based perspective, enhancing the model’s perception of target data distributions without requiring access to training data or processes. Building on this perspective, we introduce Test-time Energy Adaptation (TEA), which transforms the trained classifier into an energy-based model and aligns the model’s distribution with the test data’s, enhancing its ability to perceive test distributions and thus improving overall generalizability. Extensive experiments across multiple tasks, benchmarks and architectures demonstrate TEA’s superior generalization performance against state-of-the-art methods. Further in-depth analyses reveal that TEA can equip the model with a comprehensive perception of test distribution, ultimately paving the way toward improved generalization and calibration.
@@ 9 Limitation and Future Works
Method Noise Blur Weather Digital Avg Gaussian Shot Impulse Defocus Glass Motion Zoom Snow Frost Fog Bright Contrast Elastic Pixel JPEG Acc($\uparrow$) mCE($\downarrow$) CIFAR-10(-C) Source 27.68 34.25 27.07 53.01 45.67 65.24 57.99 74.87 58.68 73.98 90.70 53.37 73.39 41.56 69.71 56.47 100.00 BN 71.93 73.88 63.74 87.19 64.72 85.83 87.89 82.73 82.61 84.75 91.61 87.35 76.25 80.33 72.70 79.56 52.65 DUA PL 14.46 17.77 14.48 44.27 31.29 66.70 60.49 76.46 55.01 79.35 91.05 43.49 75.36 31.61 69.63 51.42 106.98 SHOT 62.91 66.15 46.41 82.85 61.64 82.39 83.16 79.03 78.24 82.35 89.68 82.20 76.08 75.34 73.19 74.77 63.19 TENT 75.20 76.52 67.03 88.00 68.22 86.27 89.24 84.06 83.81 86.30 92.16 87.91 78.01 82.77 75.77 81.41 48.13 ETA 72.21 73.88 63.72 87.21 64.70 85.84 87.89 82.73 82.58 84.77 91.60 87.32 76.22 80.37 72.69 79.58 52.64 EATA 72.25 73.88 63.74 87.19 64.72 85.83 87.89 82.73 82.61 84.75 91.61 87.35 76.25 80.33 72.70 79.59 52.62 SAR 71.95 74.14 64.11 87.39 65.20 86.00 88.06 83.08 82.66 85.07 91.90 87.20 76.69 80.41 72.79 79.77 51.94 TEA 78.33 79.87 70.94 88.89 71.31 87.87 89.77 85.56 85.29 87.61 92.37 88.98 79.32 84.90 78.99 83.33 43.69 CIFAR-100(-C) Source 9.87 11.58 4.15 35.57 20.56 46.92 44.20 51.13 39.07 45.07 71.42 27.43 51.01 30.19 42.82 35.39 100.00 BN 46.53 48.62 37.15 70.94 47.36 69.04 71.25 63.00 62.96 66.08 75.89 71.31 58.79 64.56 47.46 60.06 63.54 PL 29.54 34.09 14.99 64.06 40.81 65.36 67.74 60.42 59.47 62.92 75.98 57.29 58.72 59.41 50.34 53.40 72.12 SHOT 37.50 39.68 21.27 67.90 45.52 67.42 69.86 61.42 60.75 65.01 75.12 63.96 58.96 62.54 51.09 56.53 68.01 TENT 53.44 54.12 45.53 71.69 51.17 71.54 71.63 64.88 65.30 68.41 75.14 73.59 59.25 66.81 53.93 63.09 59.42 ETA 48.64 50.95 38.05 69.66 47.52 67.76 70.24 62.51 61.73 66.03 73.40 71.15 56.93 64.40 48.34 59.82 64.52 EATA 48.76 51.60 39.47 69.29 47.49 68.13 70.68 62.94 62.53 65.14 74.48 71.61 57.53 64.24 49.67 60.24 63.75 SAR 51.34 54.08 44.62 72.24 50.10 71.06 72.43 64.96 65.35 68.40 76.23 73.95 60.04 67.26 52.29 62.95 59.37 TEA 54.29 56.55 48.59 72.96 53.78 72.63 74.20 67.78 67.14 69.98 76.74 75.71 62.18 68.65 55.32 65.10 56.18 Tiny-ImageNet(-C) Source 12.37 16.07 5.55 6.63 5.67 20.71 19.38 26.79 31.88 14.18 29.75 1.95 32.03 49.00 46.19 21.21 100.00 BN 24.00 25.82 18.22 28.43 18.46 35.20 33.41 27.59 30.26 25.58 32.58 6.04 34.12 40.81 35.72 27.74 93.42 PL 20.75 25.99 11.41 17.28 14.42 36.07 33.66 30.19 35.01 22.44 34.87 2.03 41.31 51.76 46.75 28.26 91.22 SHOT 22.25 27.10 14.54 18.74 15.17 36.78 34.63 30.95 35.68 23.52 36.01 1.91 41.01 51.99 46.85 29.14 90.16 TENT 23.80 25.29 18.68 27.05 17.54 33.21 31.36 26.21 28.56 24.76 30.08 5.68 31.37 37.77 33.33 26.31 95.52 ETA 24.44 25.76 18.00 28.05 18.25 34.59 32.85 27.11 29.87 25.42 31.86 5.93 32.84 39.55 34.82 27.28 94.12 EATA 24.48 25.55 17.95 27.68 18.41 34.83 32.84 27.03 29.67 25.39 31.94 5.87 32.71 39.87 35.09 27.28 94.09 SAR 25.10 26.33 18.87 28.80 18.44 35.59 33.87 28.31 30.62 26.32 33.31 6.30 34.24 41.16 35.97 28.21 92.82 TEA 25.50 29.35 20.58 32.76 21.14 41.72 39.06 31.43 34.19 29.02 36.83 6.89 38.94 46.38 41.31 31.67 87.99

### 27 Feb 2024 v2

@@ Abstract
Test-time adaptation (TTA) aims to improve model generalizability when test data diverges from training distribution, offering the distinct advantage of not requiring access to training data and processes, especially valuable in the context of large pre-trained models. However, current TTA methods fail to address the fundamental issue: covariate shift, i.e., the decreased generalizability can be attributed to the model’s reliance on the marginal distribution of the training data, which may impair model calibration and introduce confirmation bias. To address this, we propose a novel energy-based perspective, enhancing the model’s perception of target data distributions without requiring access to training data or processes. Building on this perspective, we introduce Test-time Energy Adaptation (TEA), which transforms the trained classifier into an energy-based model and aligns the model’s distribution with the test data’s, enhancing its ability to perceive test distributions and thus improving overall generalizability. Extensive experiments across multiple tasks, benchmarks and architectures demonstrate TEA’s superior generalization performance against state-of-the-art methods. Further in-depth analyses reveal that TEA can equip the model with a comprehensive perception of test distribution, ultimately paving the way toward improved generalization and calibration [^1] Code is available: https://github.com/yuanyige/tea.
@@ 9 Limitation and Future Works
Method Noise Blur Weather Digital Avg Gaussian Shot Impulse Defocus Glass Motion Zoom Snow Frost Fog Bright Contrast Elastic Pixel JPEG Acc($\uparrow$) mCE($\downarrow$) CIFAR-10(-C) Source 27.68 34.25 27.07 53.01 45.67 65.24 57.99 74.87 58.68 73.98 90.70 53.37 73.39 41.56 69.71 56.47 100.00 BN 71.93 73.88 63.74 87.19 64.72 85.83 87.89 82.73 82.61 84.75 91.61 87.35 76.25 80.33 72.70 79.56 52.65 DUA* 72.60 75.40 64.70 86.90 65.10 85.40 88.40 83.20 82.50 86.90 92.40 85.90 77.30 80.70 73.80 80.10 50.78 PL 14.46 17.77 14.48 44.27 31.29 66.70 60.49 76.46 55.01 79.35 91.05 43.49 75.36 31.61 69.63 51.42 106.98 SHOT 62.91 66.15 46.41 82.85 61.64 82.39 83.16 79.03 78.24 82.35 89.68 82.20 76.08 75.34 73.19 74.77 63.19 TENT 75.20 76.52 67.03 88.00 68.22 86.27 89.24 84.06 83.81 86.30 92.16 87.91 78.01 82.77 75.77 81.41 48.13 ETA 72.21 73.88 63.72 87.21 64.70 85.84 87.89 82.73 82.58 84.77 91.60 87.32 76.22 80.37 72.69 79.58 52.64 EATA 72.25 73.88 63.74 87.19 64.72 85.83 87.89 82.73 82.61 84.75 91.61 87.35 76.25 80.33 72.70 79.59 52.62 SAR 71.95 74.14 64.11 87.39 65.20 86.00 88.06 83.08 82.66 85.07 91.90 87.20 76.69 80.41 72.79 79.77 51.94 TEA 78.33 79.87 70.94 88.89 71.31 87.87 89.77 85.56 85.29 87.61 92.37 88.98 79.32 84.90 78.99 83.33 43.69 CIFAR-100(-C) Source 9.87 11.58 4.15 35.57 20.56 46.92 44.20 51.13 39.07 45.07 71.42 27.43 51.01 30.19 42.82 35.39 100.00 BN 46.53 48.62 37.15 70.94 47.36 69.04 71.25 63.00 62.96 66.08 75.89 71.31 58.79 64.56 47.46 60.06 63.54 PL 29.54 34.09 14.99 64.06 40.81 65.36 67.74 60.42 59.47 62.92 75.98 57.29 58.72 59.41 50.34 53.40 72.12 SHOT 37.50 39.68 21.27 67.90 45.52 67.42 69.86 61.42 60.75 65.01 75.12 63.96 58.96 62.54 51.09 56.53 68.01 TENT 53.44 54.12 45.53 71.69 51.17 71.54 71.63 64.88 65.30 68.41 75.14 73.59 59.25 66.81 53.93 63.09 59.42 ETA 48.64 50.95 38.05 69.66 47.52 67.76 70.24 62.51 61.73 66.03 73.40 71.15 56.93 64.40 48.34 59.82 64.52 EATA 48.76 51.60 39.47 69.29 47.49 68.13 70.68 62.94 62.53 65.14 74.48 71.61 57.53 64.24 49.67 60.24 63.75 SAR 51.34 54.08 44.62 72.24 50.10 71.06 72.43 64.96 65.35 68.40 76.23 73.95 60.04 67.26 52.29 62.95 59.37 TEA 54.29 56.55 48.59 72.96 53.78 72.63 74.20 67.78 67.14 69.98 76.74 75.71 62.18 68.65 55.32 65.10 56.18 Tiny-ImageNet(-C) Source 12.37 16.07 5.55 6.63 5.67 20.71 19.38 26.79 31.88 14.18 29.75 1.95 32.03 49.00 46.19 21.21 100.00 BN 24.00 25.82 18.22 28.43 18.46 35.20 33.41 27.59 30.26 25.58 32.58 6.04 34.12 40.81 35.72 27.74 93.42 PL 20.75 25.99 11.41 17.28 14.42 36.07 33.66 30.19 35.01 22.44 34.87 2.03 41.31 51.76 46.75 28.26 91.22 SHOT 22.25 27.10 14.54 18.74 15.17 36.78 34.63 30.95 35.68 23.52 36.01 1.91 41.01 51.99 46.85 29.14 90.16 TENT 23.80 25.29 18.68 27.05 17.54 33.21 31.36 26.21 28.56 24.76 30.08 5.68 31.37 37.77 33.33 26.31 95.52 ETA 24.44 25.76 18.00 28.05 18.25 34.59 32.85 27.11 29.87 25.42 31.86 5.93 32.84 39.55 34.82 27.28 94.12 EATA 24.48 25.55 17.95 27.68 18.41 34.83 32.84 27.03 29.67 25.39 31.94 5.87 32.71 39.87 35.09 27.28 94.09 SAR 25.10 26.33 18.87 28.80 18.44 35.59 33.87 28.31 30.62 26.32 33.31 6.30 34.24 41.16 35.97 28.21 92.82 TEA 25.50 29.35 20.58 32.76 21.14 41.72 39.06 31.43 34.19 29.02 36.83 6.89 38.94 46.38 41.31 31.67 87.99